# On Sensitivity of Learning with Limited Labelled Data to the Effects of Randomness: Impact of Interactions and Systematic Choices

### v1

@@ Appendix C Ablation Study: Reducing Number of Mitigation Runs and Test Data Size
As mentioned in Section 3, there is a trade-off between feasibility (computations costs) of the investigation and precision (reliability) of the investigation results. This trade-off mainly depends on the number of mitigation runs (i.e., the number of configurations explored for the non-investigated randomness factors). To determine the optimal number of mitigation runs, we explore the effects of reducing the number of mitigation runs ($M$) and the number of test samples used for evaluation on the reliability of the results. We perform this ablation study for the Flan-T5 model on the SST2 dataset.
@@
Even when reducing the computation cost of the proposed method to $\sim$1of our baseline setting (reducing the number of mitigation runs to 10 and using only 1 000 test samples for evaluation) the findings can be considered sufficiently reliable. Therefore, if the precision of the results is not as paramount, the proposed method can be used even in this reduced setting (although one needs to be aware of the implications). To produce more precise results, and due to the significant computation costs of running the larger in-context learning models (LLaMA-2, Mistral and Zephyr), we have decided to investigate the stability of these models using 20 mitigation runs and 1 000 test samples (following the practice in related work Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2023); Li and Qiu (2023); Köksal et al. (2023)). As such, the observed importance scores for different factors may be affected by this choice, but the findings regarding the importance should still hold. In addition, to keep the comparison between models as unbiased as possible, we use the same amount of test data for all the models and all the datasets and across all experiments.
@@ Appendix D Experimental Setup and Implementation Details
SST-2 A
{Negative, Positive}
Determine sentiment of the sentence using following options: 1) [Class 1] 2) [Class 2].
@@
B
Same as above
[Input] Sentiment? [Output]
C
Same as above
[Input] Sentiment is [Output]
D
{terrible, great}
[Input] It was [Output]
CoLA A
{No, Yes}
Determine grammatical acceptability of the sentence using following options: 1) [Class 1] 2) [Class 2].
@@
B
Same as above
[Input] Grammatically acceptable? [Output]
C
{Yes, No}
[Input] Grammar problems? [Output]
D
{not acceptable, acceptable}
[Input] It is [Output]
MRPC A
{No, Yes}
Determine whether the sentence pair is semantically equivalent using following options: 1) [Class 1] 2) [Class 2].
@@
B
Same as above
[Input] Semantically equivalent sentences? [Output]
C
{Yes, No}
[Input] Semantically different sentences? [Output]
D
{not equivalent, equivalent}
[Input] Sentences are [Output]
AG News A
{World, Sports, Business, Science and Technology}
Determine topic of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@
B
Same as above
[Input] Topic? [Output]
C
Same as above
[Input] Topic is [Output]
D
Same as above
User: [Input] This is about [Output]
TREC A
{Expression, Entity, Description, Human, Location, Number}
Determine topic of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@
B
Same as above
[Input] Topic? [Output]
C
Same as above
[Input] Topic is [Output]
D
Same as above
User: [Input] This is about [Output]
SNIPS A
{Playlist, Weather, Event, Musing, Creative Work, Rate Book, Book Restaurant}
Determine intent of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@
B
Same as above
[Input] Intent? [Output]
C
Same as above
[Input] Intent is [Output]
D
Same as above
User: [Input] User requested [Output]
DB-Pedia A
{Company, Educational Institution, Artist, Athlete, Office Holder, Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, Written Work}
Determine topic of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@
B
Same as above
[Input] Topic? [Output]
C
Same as above
[Input] Topic is [Output]
D
Same as above
User: [Input] This is about [Output]
@@
Based on the ablation study (included in Appendix 3), we use 10 investigation and 20 mitigation runs (resulting in overall 200 training and evaluation runs) for the larger models (LLaMA-2, Mistral-7B and Zephyr-7B) and 100 mitigation runs (results in overall 1 000 training and evaluation runs) for the smaller models (Flan-T5, BERT, RoBERTa). Following the practice from the related work (e.g., Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2023); Li and Qiu (2023); Köksal et al. (2023)) and the results of our ablation study, we evaluate each run using only 1 000 test samples (the selection is governed by the Label Selection randomness factor). The main reason is the computation cost of the inference for the large language models. However, to keep the comparisons between all the models as unbiased as possible, we use the same amount of test samples for the transfer learning and meta-learning approaches as well. These decisions represents the trade-off between the precision/reliability of the results and their feasibility/required computation costs.
@@
For the fine-tuning models, BERT [^6] https://huggingface.co/bert-base-uncased and RoBERTa [^7] https://huggingface.co/roberta-base, we use the base version of the pre-trained models from HuggingFace Wolf et al. (2019). Both models are trained in full (without freezing the pre-trained part) on all datasets using learning rate of 1e-5 for 5 epochs on binary and 10 epochs on multi-class dataset, using early stopping, Adam optimiser and batch size 8.
@@
Appendix E Additional Results: Impact of Interactions Between Randomness Factors and Comparison with Typical Investigation Methods
@@
To showcase the impact of interactions between randomness factors on the investigation of the effects of the different randomness factors, we provide a comparison between typical investigation strategies from related work and our proposed method:
@@
Flan-T5 Random Fixed Interactions
Golden Model 2.244 2.244 2.244
Label Select. (*) 2.517 (*) 2.594 (*) 2.128
Data Split (*) 2.362 (*) 2.480 (*) 2.167
Data Order (*) 2.131 (*) 3.014 0.869
Sample Choice (*) 2.370 (*) 3.191 (*) 2.123
Zephyr-7B Random Fixed Interactions
Golden Model 1.043 1.043 1.043
Label Select. (*) 1.122 (*) 1.004 (*) 0.863
Data Split (*) 1.185 0.402 (*) 0.664
Data Order (*) 1.138 (*) 0.957 0.456
Sample Choice (*) 1.052 0.406 (*) 0.744
BERT Random Fixed Interactions
Golden Model 0.970 0.970 0.970
Label Select. (*) 1.096 (*) 1.409 (*) 0.927
Data Split (*) 1.096 (*) 1.272 (*) 0.937
Model Init. (*) 1.155 (*) 1.197 (*) 0.828
Data Order (*) 1.082 (*) 1.217 (*) 0.852
ProtoNets Random Fixed Interactions
Golden Model 0.940 0.940 0.940
Label Select. (*) 0.987 (*) 0.857 (*) 0.887
Data Split (*) 1.012 (*) 1.041 (*) 0.959
Model Init. (*) 0.892 (*) 0.845 0.658
Data Order (*) 0.929 (*) 3.510 (*) 3.233
Sample Choice (*) 0.983 (*) 0.832 0.646
Table 5: Comparison of different investigation strategies for the Flan-T5, Zephyr-7B, BERT fine-tuning and meta-learning with the Prototypical Networks on the SST2 dataset. The ‘Random‘ strategy simply repeats the training and evaluation multiple times without any constraints. In the ‘Fixed‘ strategy, the randomness factor configuration is kept fixed to a single state during investigation. We compare these investigation strategies with our proposed method. We run each investigation strategy the same number of times (denoted as ‘Full‘, number of runs is governed by our method). Our method (‘Interactions‘) takes the interactions into consideration. Factors considered important for different strategies are denoted using the (*) symbol. We observe that interactions between factors may cause some factors to have their importance overestimated (denoted in bold) or underestimated (denoted in italics).
To keep the comparison between the investigation strategies as unbiased as possible, we perform same number of training and evaluation runs for each one. For each strategy, we repeat the training and evaluation 1 000 times (or 200 times, as governed by the number of runs in our proposed method). For the ‘Random‘ and ‘Fixed‘ strategy, we consider factor to be important when it contributes at least 50% of the golden model standard deviation. The importance of factors in our method is determined using the importance score. The full results are presented in Table 5.
Looking at the results of the ‘Random‘ investigation strategy, we can see that the deviation contributed by the different randomness factors is, more or less, consistent across all the investigated randomness factors. At the same time, this contributed deviation is similar to the one from golden model. As such, we observe over-estimation of some specific randomness factors. Even though using the ‘Fixed‘ investigation strategy produces more reliable results (which are more distributed and start to follow the results as found by our proposed method), it is still affected by the randomness caused by the choice of the single randomness factor configurations for the non-investigated factors. The results still show both over-estimation and under-estimation of effects for the randomness factors.
Overall using either ‘Random‘ or ‘Fixed‘ investigation strategy appears to lead to incorrect findings and incorrect attribution of the effects of different randomness factors (either over-estimating impact of non-important randomness factors or under-estimating the impact of the important factors). For example, in the case of Flan-T5 in-context learning, these investigation strategies indicate that all the randomness factors are equally important (as they contribute similar deviation to the golden model), which is not the case when the interactions are taken into consideration (when interactions are considered, the impact of data order falls off). From the results, we can see over-estimation for the effects of randomness factors that are not considered important by our proposed method, while at the same time we see under-estimation for some of the randomness factors that are considered important. Therefore, the use of these strategies may lead to inconsistency of results, as it is highly dependent on the underlying randomness (in ‘Fixed‘ strategy) or the effects are not correctly attributed to the specific randomness factors (as is the case with ‘Random‘ strategy).
Finally, these strategies limit the possibilities for analysing the importance of different randomness factors. As only a single deviation value is produced for each factor, only importance relative to the golden model can be determined based on a threshold value. However, determining such heuristic may be problematic in some cases, as the deviation from different randomness factors are often very close to each other when using these strategies (especially for the ‘Random‘ investigation strategy). In addition, the relative importance of the factors can be determined by ordering these factors based on their contributed deviation. However, this relative ordering is not a good indicator as it introduces further dependence on randomness (in case of both strategies, we see at least 4/5 unique ordering of the randomness factors). As such, we cannot determine the importance with a good reliability. As our method can determine this importance, it can be considered superior in this regard as well.

### v2

@@ Appendix C Ablation Study: Reducing Number of Mitigation Runs and Test Data Size
As mentioned in Section 3, there is a trade-off between feasibility (computations costs) of the investigation and precision (reliability) of the investigation results. This trade-off mainly depends on the number of mitigation runs (i.e., the number of configurations explored for the non-investigated randomness factors). To determine the optimal number of mitigation runs, we explore this trade-off using a modified version of the method described in Appendix B.3: we run the investigation for a larger number of mitigation runs (observing the behaviour even after the optimal point) and explore the effects of reducing the number of mitigation runs ($M$) and the number of test samples used for evaluation on the results and how well they estimate the overall contributed effects and how well the interactions are mitigated. We perform this ablation study for the Flan-T5 model on the SST2 dataset and report only specific interesting points.
@@
Even when reducing the computation cost of the proposed method to $\sim$1of our baseline setting (reducing the number of mitigation runs to 10 and using only 1 000 test samples for evaluation) the findings can be considered sufficiently reliable. Therefore, if the precision of the results is not as paramount, the proposed method can be used even in this reduced setting (although one needs to be aware of the implications). To produce more precise results, and due to the significant computation costs of running the larger in-context learning models (LLaMA-2, Mistral and Zephyr), we have decided to run the investigation using 20 mitigation runs and 1 000 test samples (following the practice in related work Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2023); Li and Qiu (2023); Köksal et al. (2023)). As such, the observed importance scores for different factors may be affected by this choice, but the findings regarding the importance should still hold. In addition, to keep the comparison between models as unbiased as possible, we use the same amount of test data for all the models and all the datasets and across all experiments.
@@ Appendix D Experimental Setup and Implementation Details
SST-2 A {Negative, Positive} Determine sentiment of the sentence using following options: 1) [Class 1] 2) [Class 2].
@@
B Same as above [Input] Sentiment? [Output]
C Same as above [Input] Sentiment is [Output]
D {terrible, great} [Input] It was [Output]
CoLA A {No, Yes} Determine grammatical acceptability of the sentence using following options: 1) [Class 1] 2) [Class 2].
@@ C Same as above [Input] Sentiment is [Output]
B Same as above [Input] Grammatically acceptable? [Output]
C {Yes, No} [Input] Grammar problems? [Output]
D {not acceptable, acceptable} [Input] It is [Output]
MRPC A {No, Yes} Determine whether the sentence pair is semantically equivalent using following options: 1) [Class 1] 2) [Class 2].
@@ B Same as above [Input] Grammatically acceptable? [Output]
B Same as above [Input] Semantically equivalent sentences? [Output]
C {Yes, No} [Input] Semantically different sentences? [Output]
D {not equivalent, equivalent} [Input] Sentences are [Output]
AG News A {World, Sports, Business, Science and Technology} Determine topic of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@ B Same as above [Input] Semantically equivalent sentences? [Output]
B Same as above [Input] Topic? [Output]
C Same as above [Input] Topic is [Output]
D Same as above User: [Input] This is about [Output]
TREC A {Expression, Entity, Description, Human, Location, Number} Determine topic of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@ D Same as above User: [Input] This is about [Output]
B Same as above [Input] Topic? [Output]
C Same as above [Input] Topic is [Output]
D Same as above User: [Input] This is about [Output]
SNIPS A {Playlist, Weather, Event, Musing, Creative Work, Rate Book, Book Restaurant} Determine intent of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@
B Same as above [Input] Intent? [Output]
C Same as above [Input] Intent is [Output]
D Same as above User: [Input] User requested [Output]
DB-Pedia A {Company, Educational Institution, Artist, Athlete, Office Holder, Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, Written Work} Determine topic of the sentence using following options: 1) [Class 1] 2) [Class 2] … N) [Class N].
@@ D Same as above User: [Input] User requested [Output]
B Same as above [Input] Topic? [Output]
C Same as above [Input] Topic is [Output]
D Same as above User: [Input] This is about [Output]
@@
Based on the ablation study (included in Appendix 3), we use 10 investigation and 20 mitigation runs (resulting in overall 200 training and evaluation runs) for the in-context learning (Flan-T5, LLaMA-2, Mistral-7B and Zephyr-7B) and 100 mitigation runs (results in overall 1 000 training and evaluation runs) for the other approaches that use smaller models (BERT, RoBERTa). Following the practice from the related work (e.g., Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2023); Li and Qiu (2023); Köksal et al. (2023)) and the results of our ablation study, we evaluate each run using only 1 000 test samples (the selection is governed by the Label Selection randomness factor). The main reason is the computation cost of the inference for the large language models. To prevent introduction of any biases into the comparison, we use the same amount of test samples for the transfer learning and meta-learning as well (although we use larger number of runs results in larger distributions in those cases). These decisions represents the trade-off between feasibility/required computation costs to achieve the results and how well the effects of randomness factors are estimated and the interactions between them mitigated.
@@
For the fine-tuning models, BERT [^6] https://huggingface.co/bert-base-uncased and RoBERTa [^7] https://huggingface.co/roberta-base, we use the base version of the pre-trained models from HuggingFace Wolf et al. (2019). Both models are trained in full (without freezing the pre-trained part) on all datasets using learning rate of 1e-5 for 5 epochs on binary and 10 epochs on multi-class dataset, using early stopping, AdamW optimiser with warmup for 10% of the steps and batch size 8.
@@
Appendix E Validating the Proposed Investigation Method
In this Appendix, we provide further information on how the proposed investigation method was validated. As there is no ground-truth of the effects of randomness to compare against, the validity of methods for investigating the effects of randomness can be evaluated only indirectly. In this paper, we perform such indirect evaluation/validation by:
  1. 1.
    Evaluating the properties and benefits of the proposed methods by comparing it to the existing ones. As discussed in the main content of the paper, the benefits of the proposed methods are: 1) importance score that can be used for more in-depth analysis (relative ordering of randomness factors and comparison across models, datasets and other experimental settings), as opposed to determining the importance only in binary fashion from previous works (factor is or is not important); and 2) handling interactions between effects of randomness factors, which, when previously ignored or not being addressed sufficiently caused inconsistencies in findings. We discuss this validation further in Appendix E.1 (which we consider as the main validation of the method).
  2. 2.
    Exploring how the results and findings change as we change the overall number of runs. The results and findings of the method are dependent on the choice of how many investigation and mitigation runs are used. We observe a trade-off between how well the results are estimated (higher number of investigation runs leads to better estimation) and the interactions mitigation (higher number of mitigation runs leads to better mitigation), and the computation costs required to achieve the results (increasing the number of runs increases the overall costs). We discuss this validation further in Appendix E.2.
  3. 3.
    Applying the method to different settings (factors, models, datasets) and observing the consistency of its results and findings. As the investigation method is designed to be general, it should be applied across different experimental settings without showing any problems (i.e., working out-of-the-box on multiple factors, models and datasets). We discuss this validation further in Appendix E.3.
E.1 Additional Results: Validation of Method Through Comparison with Typical Investigation Strategies
SST2 CoLA MRPC
Flan-T5 Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 2.244 2.244 2.244 3.811 3.811 3.811 1.328 1.328 1.328
Label Select. (*) 2.517 (*) 2.594 (*) 2.128 (*) 3.602 (*) 2.804 (*) 3.257 (*) 1.122 (*) 1.189 0.363
Data Split (*) 2.362 (*) 2.480 (*) 2.167 (*) 3.961 (*) 1.990 (*) 3.483 (*) 1.503 0.252 (*) 0.926
Data Order (*) 2.131 (*) 3.014 0.869 (*) 3.122 (*) 4.172 1.793 (*) 1.007 0.289 0.209
Sample Choice (*) 2.370 (*) 3.191 (*) 2.123 (*) 3.478 1.203 (*) 3.138 (*) 1.277 (*) 0.678 0.348
Zephyr-7B Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 1.043 1.043 1.043 9.566 9.566 9.566 12.785 12.785 12.785
Label Select. (*) 1.122 (*) 1.004 (*) 0.863 (*) 7.367 2.529 (*) 5.806 (*) 11.968 (*) 11.977 5.109
Data Split (*) 1.185 0.402 (*) 0.664 (*) 8.235 (*) 8.001 (*) 7.675 (*) 12.504 (*) 6.660 5.973
Data Order (*) 1.138 (*) 0.957 0.456 (*) 9.622 (*) 7.028 3.598 (*) 11.211 4.913 (*) 8.038
Sample Choice (*) 1.052 0.406 (*) 0.744 (*) 10.069 4.379 (*) 9.135 (*) 12.980 (*) 12.239 6.305
BERT Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 0.970 0.970 0.970 1.473 1.473 1.473 2.929 2.929 2.929
Label Select. (*) 1.096 (*) 1.409 (*) 0.927 (*) 1.552 (*) 1.103 (*) 1.212 (*) 2.760 (*) 2.308 (*) 2.168
Data Split (*) 1.096 (*) 1.272 (*) 0.937 (*) 1.409 (*) 1.649 (*) 1.250 (*) 2.904 (*) 3.132 (*) 2.384
Model Init. (*) 1.155 (*) 1.197 (*) 0.828 (*) 1.523 (*) 2.481 (*) 1.059 (*) 2.813 (*) 1.997 (*) 2.180
Data Order (*) 1.082 (*) 1.217 (*) 0.852 (*) 1.639 (*) 1.333 1.086 (*) 2.809 (*) 3.971 (*) 2.081
ProtoNets Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 0.940 0.940 0.940 2.111 2.111 2.111 1.789 1.789 1.789
Label Select. (*) 0.987 (*) 0.857 (*) 0.887 (*) 2.109 (*) 1.497 (*) 1.924 (*) 1.572 0.451 (*) 1.448
Data Split (*) 1.012 (*) 1.041 (*) 0.959 (*) 2.188 (*) 2.301 (*) 2.010 (*) 1.919 (*) 1.006 (*) 1.791
Model Init. (*) 0.892 (*) 0.845 0.658 (*) 2.222 (*) 1.582 (*) 1.801 (*) 1.888 0.610 1.240
Data Order (*) 0.929 (*) 3.510 (*) 3.233 (*) 4.114 0.439 (*) 3.346 (*) 3.087 (*) 6.590 (*) 2.265
Sample Choice (*) 0.983 (*) 0.832 0.646 (*) 2.163 0.890 (*) 1.659 (*) 1.805 (*) 1.271 1.084
AG News TREC SNIPS
Flan-T5 Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 3.090 3.090 3.090 1.324 1.324 1.324 2.284 2.284 2.284
Label Select. 0.980 0.391 0.556 (*) 1.502 (*) 1.210 0.683 (*) 3.081 (*) 2.969 1.581
Data Split (*) 6.152 0.594 (*) 3.777 (*) 1.247 (*) 1.844 0.892 (*) 2.855 (*) 2.740 1.602
Data Order (*) 2.962 (*) 4.912 0.686 (*) 1.222 (*) 1.005 (*) 0.815 (*) 2.040 0.964 (*) 1.769
Sample Choice (*) 1.982 1.466 (*) 0.806 (*) 1.616 (*) 1.344 0.819 (*) 3.156 0.970 1.590
Zephyr-7B Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 2.066 2.066 2.066 3.884 3.884 3.884 4.132 4.132 4.132
Label Select. (*) 1.966 1.008 (*) 1.460 (*) 3.196 (*) 3.988 (*) 2.963 (*) 2.687 (*) 2.812 (*) 2.935
Data Split (*) 2.141 (*) 2.452 (*) 1.817 (*) 3.554 (*) 2.115 (*) 3.221 (*) 4.006 (*) 3.580 (*) 3.052
Data Order (*) 1.859 (*) 2.243 0.919 (*) 4.037 (*) 3.990 1.925 (*) 3.838 1.012 (*) 3.007
Sample Choice (*) 2.358 0.884 (*) 1.874 (*) 4.021 (*) 4.696 (*) 3.279 (*) 3.975 1.311 (*) 3.331
BERT Random Fixed Interactions Random Fixed Interactions Random Fixed Interactions
Golden Model 1.239 1.239 1.239 1.667 1.667 1.667 0.486 0.486 0.486
Label Select. (*) 1.202 (*) 0.923 (*) 0.979 (*) 1.600 (*) 1.513 (*) 1.348 (*) 0.559 (*) 0.405 (*) 0.308
Data Split (*) 1.462 (*) 1.365 (*) 1.164 (*) 1.502 (*) 1.513 (*) 1.568 (*) 0.401 (*) 0.426 (*) 0.294
Model Init. (*) 1.142 (*) 1.047 0.693 (*) 1.926 (*) 1.108 0.939 (*) 0.479 (*) 0.635 0.121
Data Order (*) 1.335 (*) 0.714 0.686 (*) 1.666 (*) 1.391 1.019 (*) 0.471 0.173 0.103
Table 5: Comparison of different investigation strategies for the Flan-T5, Zephyr-7B and BERT fine-tuning on the binary datasets (SST2, CoLA and MRPC) and the multi-class datasets (AG News, TREC and SNIPS). Comparison for the DB Pedia dataset is not included as Flan-T5 model shows poor performance on this particular dataset. The ‘Random‘ strategy simply repeats the training and evaluation multiple times without any constraints. In the ‘Fixed‘ strategy, the randomness factor configuration is kept fixed to a single state during investigation. We compare these investigation strategies with our proposed method. We run each investigation strategy the same number of times (number of runs is governed by our method). Our method (‘Interactions‘) takes the interactions into consideration. Factors considered important for different strategies are denoted using the (*) symbol. We observe that interactions between factors may cause some factors to have their importance overestimated (denoted in bold) or underestimated (denoted in italics).
@@ E.1 Additional Results: Validation of Method Through Comparison with Typical Investigation Strategies
To showcase the impact of interactions between randomness factors on the investigation of the effects of different randomness factors, and to showcase the properties and benefits of our proposed method, we provide a comparison between the typical investigation strategies from related work and our proposed method:
@@
To prevent introduction of any biases into the comparisons between the strategies, we perform same number of training and evaluation runs for each method. For each strategy, we repeat the training and evaluation 1 000 times (or 200 times, as governed by the number of runs in our proposed method). The full results are presented in Table 5 (except for DB Pedia dataset where the Flan-T5 model does not work well). We focus on 2 main aspects in the comparison: 1) determining importance of the factors; 2) how interactions affect the findings.
Determining importance of the factors.
As the Random and Fixed strategies results only in a single score (deviation in the results), we consider the factor to be important when it contributes at least 50% of the golden model standard deviation. As such, the importance of the randomness factors can be determined only in a binary fashion (factor is or is not important). Such setting allows only for a limited analysis (only relative ordering of factors based on the deviation withing the same setup) and cannot be easily used to compare the importance across different models. On the other hand, our proposed method provides an importance score that can be used for more in-depth analysis, such as the relative ordering of randomness factors based on their importance, or comparison across models, datasets and experimental settings (as the importance score is normalised with the overall deviation in the results from the golden model). This benefit can be illustrated using following example – using Table 5 and the Random/Fixed strategy, we cannot say with good conscience that the Sample Choice is more important for the Flan-T5 model than for the Zephyr-7B model based only on their standard deviation ($2.370$ vs. $1.052$ using Random; $3.191$ vs. $0.406$ using Fixed) as the overall deviation in results is higher, but can be done so using our method (importance score from Figure 1 or Table 6 and 9 of $0.57$ vs. $0.39$) as the score is normalised. Or similarly for Data order ($2.131$ vs. $1.138$ for Random; $3.014$ vs. $0.957$ for Fixed; $-0.47$ vs. $-0.44$ for our method) -– using the importance score we see the importance of Data Order is similar (slightly higher for Zephyr-7B) for both models, while other investigation strategies show a large difference (higher importance for Flan-T5).
Handling interactions.
The existing strategies either ignore the interactions completely (Random) or do not addresses the sufficiently (i.e., in a way that strongly depends on randomness in the Fixed strategy). As such, the baseline strategies often lead to incorrect attribution of the effects of different factors, either due to overestimating the impact of non-important randomness factors, or underestimating the impact of important factors. For example, in the case of Flan-T5 in-context learning, these investigation strategies indicate that all the randomness factors are equally important (as they contribute similar deviation to the golden model), which is not the case when the interactions are taken into consideration (when interactions are considered, the impact of data order falls off). In case of the Random strategy, this behaviour stems from the strategy consistently leading to the same overall deviation/importance for all the investigated randomness factors (which is similar to the deviation of the Golden Model). Even though using the Fixed investigation strategy produces more reliable results (which are more distributed and handle the interactions to a certain extent), it is still affected by the randomness caused by the choice of the single randomness factor configurations for the non-investigated factors. The results still show both overestimation and underestimation of effects for the randomness factors. On the other hand, our method is specifically designed to handle the interactions using the mitigation runs. Handling the interactions this way, we observe that the finding that the long believed sensitivity of in-context learning to Data Order is actually a sensitivity to Sample Choice (and potentially the choice of prompt format) when choosing samples in a more sophisticated manner, holds even when choosing samples at random.
All in all, our proposed method provides 2 significant benefits over the baseline strategies, which indirectly validates its use: 1) allowing for more in-depth analysis and comparison across different factors, models, datasets and experimental setups that leads to actionable findings and read-to-apply take-away messages and suggestions (described in experimental results in Sections 4.2, 4.3 and 4.4, such as increasing the number of shots for in-context learning reduces the importance of sample choice, but does not affect the importance of sample order); and 2) handling of interactions that leads to more consistent results.
E.2 Additional Results: Validation of Method by Exploring the Changes Due to Different Number of Runs
The results and findings from investigation are heavily dependent on the overall number of runs. As opposed to the baseline strategies, our proposed method introduces another parameter, number of mitigation runs, to handle the interactions. We provide results from exploring how changing the number of investigation and mitigation runs affect the results and findings (i.e., how well the effects are estimated and the interactions mitigated) in Appendix B.3 and Appendix C, while in this section, we provide a summary of relevant results.
The effects of randomness factors can be estimated using a relatively low number of investigation runs (around $6$ to $8$). Increasing the number of investigation runs further does not lead to considerable changes in the estimated effects (the contributed standard deviation changes only in second decimal place).
On the other hand, increasing the number of mitigation runs has a larger impact on the overall results and findings (and the different metrics we use), as it represents the main avenue for mitigating the interactions. Any change to the number of mitigation runs changes all the metrics (contributed std, mitigated std, and the importance score). In addition, the number of mitigation runs also depends on the approach, model and dataset used. As such, it is important to find the optimal point, where the interactions are sufficiently handled without requiring extensive computation costs. To find this optimal point, we provide heuristics and a simple search method in Appendix B.3. However, the overall number of required mitigation runs is still relatively low – in our experiments, we observed that using 20 mitigation runs provides sufficient mitigation of interactions and estimation of the overall effects.
Finally, we observed that the number of test samples used for evaluation is the most important factor influencing the estimation of the effects. In our experiments, we observed that using 1000 test samples for evaluation provides a good trade-off between the feasibility of larger scale experiments (due to computation costs) and the validity of the results.
E.3 Additional Results: Validation of Method by Observing Consistency of Results and Findings Across Different Settings
The proposed investigation method is designed to not be dependent on any specific experimental setup, so that it can be used across any randomness factors, model, dataset or other systematic changes (e.g., number of samples, or prompt formats). To validate this property of the method, we apply it across various settings and observe how consistent are the results and findings (the full results presented throughout the paper, such as in Tables 6-14, or Figures 1, 2, 3, or in Appendix F):
  • Different randomness factors that require different configuration setup for investigation (e.g., different choice of data, order of samples, initialisation, etc.), but also different setup for their mitigation. As discussed in Appendix B.1, the mitigation can be done on group level (effectively mitigating multiple randomness factors at the same time), while also allowing for further extensions (such as using different mitigation strategies).
  • Different approaches, namely in-context learning, fine-tuning and meta-learning, and different models in these approaches. Although each approach works differently (e.g., fine-tuning using optimisation, while in-context learning uses only inference with prompts), the proposed method works with any such approach. The only limitation is that the models and approaches used have an option to allow for deterministic behaviour. Without this option, the method can still be applied but may produce inconsistent and non-reproducible results and findings (i.e., the importance score in such cases is affected by the non-determinism of the model and so cannot be trusted fully). In addition, we apply the proposed method to models that lead to different performance and show different overall deviation in the results. In all the cases, the produced importance score can be used for the analysis and comparison, even in cases when the impact of the randomness factor is significant (e.g., Prototypical Networks on the SST2 datasets with Data Order randomness factor, where we observe a significant drop in performance and increase in overall deviation as opposed to the golden model – in which case the method correctly identifies this factor to be significantly important, leading to importance score of $0.92$)
  • Datasets with different characteristics and different experimental setups, such as different number of classes, samples, different prompt formats.
In all of these cases, the proposed method produces consistent results and findings without any obvious shortcomings. Although the baseline strategies can also be applied across all the settings, they often lead to inconsistent results due to the mishandling of interactions (i.e., Random consistently leads to results similar to Golden Model for all the randomness factors, while using Fixed strategy the importance of different factors changes quite often across different models, approaches, datasets and experimental settings).